# Sharpness-Aware Minimization and the Edge of Stability

### v5

@@ 1 Introduction
Meanwhile, Cohen et al. (2021), building on the work of Jastrzebski et al. (2019) and others, exposed a striking phenomenon regarding neural network training with the original gradient descent (GD) method: for many initialization schemes and learning rates $\eta$, the operator norm of the Hessian eventually settles in the neighborhood of $2/\eta$. This has been termed the “edge of stability”, in part because a convex quadratic trained by gradient descent with a learning rate $\eta$ will only converge if the operator norm of its Hessian (which is the same everywhere) is less than $2/\eta$. This phenomenon also inspired substantial theoretical research (see Arora et al., 2022; Damian et al., 2023; Ma et al., 2022; Zhu et al., 2022; Ahn et al., 2022; Chen and Bruna, 2022) – one result identified conditions under which, when training approaches the edge of stability, the dynamics includes a self-stabilization mechanism that tends to drive the operator norm of the Hessian back down (Damian et al., 2023).

### v6

@@ 1 Introduction
Meanwhile, Cohen et al. (2021), building on the work of Jastrzebski et al. (2019) and others, exposed a striking phenomenon regarding neural network training with the original gradient descent (GD) method: for many initialization schemes and learning rates $\eta$, the operator norm of the Hessian eventually settles in the neighborhood of $2/\eta$. This has been termed the “edge of stability”, in part because a convex quadratic trained by gradient descent with a learning rate $\eta$ will only converge if the operator norm of its Hessian (which is the same everywhere) is less than $2/\eta$. This phenomenon also inspired substantial theoretical research (see Arora et al., 2022; Damian et al., 2023; Ma et al., 2022; Zhu et al., 2022; Ahn et al., 2022; Chen and Bruna, 2022). One result identified conditions under which, when training approaches the edge of stability, the dynamics includes a self-stabilization mechanism that tends to drive the operator norm of the Hessian back down (Damian et al., 2023).